OpenStudy (anonymous):

factorize x^2 + x/4-1/8

OpenStudy (anonymous):

x^2 + x/4 - 1/8 = x^2 + x/2 - x/4 - 1/8 = x{x+ 1/2} - 1/4{ x+ 1/2} = ( x+ 1/2 ) ( x- 1/4)